OpenStudy (anonymous):

(2x-4)/(4x^2 -25) - (1)/(4x-10) = (5)/ (6x+15)

OpenStudy (anonymous):

HI :DD

OpenStudy (anonymous):

Hi. Can you help?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Can you please help??

OpenStudy (anonymous):

\[\frac{2x-4}{4x^2 -25} - \frac{1}{4x-10} = \frac{5}{6x+15}\]i just needed to write it first is that right?

OpenStudy (anonymous):

Yesss

OpenStudy (anonymous):

i was afraid so ok lets factor the denominators

OpenStudy (anonymous):

\[\frac{2x-4}{4x^2 -25} - \frac{1}{4x-10} = \frac{5}{6x+15}\] \[\frac{2x-4}{(2x+5)(2x-5)} - \frac{1}{2(x-5)} = \frac{5}{3(2x+5)}\]

OpenStudy (anonymous):

Wouldn't the second one be 2(2x-5)

OpenStudy (anonymous):

yes, you are right

OpenStudy (anonymous):

\[\frac{2x-4}{(2x+5)(2x-5)} - \frac{1}{2(2x-5)} = \frac{5}{3(2x+5)}\]

OpenStudy (anonymous):

I did that. After that I got lost

OpenStudy (anonymous):

least common multiple of the denominators is \(6(2x+5)(2x-5)\) so we can clear the fractions if we multiply both sides by it, cancelling as we go

OpenStudy (anonymous):

\[6(2x+5)(2x-5)\left(\frac{2x-4}{(2x+5)(2x-5)} - \frac{1}{2(2x-5)} \right)\]\[=6(2x+5)(2x-5)\left( \frac{5}{3(2x+5)}\right)\]

OpenStudy (anonymous):

just a lot of cancellation here you get no fractions now \[6(2x-4)-3(2x+5)=2\times 5(2x-5)\]

OpenStudy (anonymous):

now it should be a lot easier multiply out, combine like terms, etc gives \[12x-24-6x-15=20x-50\] and so on you good from there?

OpenStudy (anonymous):

Can you show me how you canceled them out???

OpenStudy (anonymous):

\[6(2x+5)(2x-5)\left(\frac{2x-4}{(2x+5)(2x-5)} - \frac{1}{2(2x-5)} \right)\] \[6(2x+5)(2x-5)\left(\frac{2x-4}{(2x+5)(2x-5)}\right) -6(2x+5)(2x-5) \left(\frac{1}{2(2x-5)}\right)\] \[6\cancel{(2x+5)(2x-5)}\left(\frac{2x-4}{\cancel{(2x+5)(2x-5)}}\right)\]\[ -\cancel{6}^3(2x+5)\cancel{(2x-5)} \left(\frac{1}{\cancel{2(2x-5)}}\right)\] \[=6(2x-4)-3(2x+5)\]

OpenStudy (anonymous):

got rid of all the common factors top and bottom

OpenStudy (anonymous):

kind of long to draw out, and that was just the left hand side

OpenStudy (anonymous):

I get it now!!!!

OpenStudy (anonymous):

great! get rid of common factors, end up with a nice linear equation to solve but you still have to be careful

OpenStudy (anonymous):

in any case if you finish solving the last equation i think you get \(\frac{11}{14}\) let me know if it works out